OpenStudy (anonymous):

3sq. rt of( x - 2) = 2sq. rt of( x)

OpenStudy (anonymous):

square both sides 9(x-2)=4x can you get it from there?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

well then it goes to 9x-18=4x -18=-5x -18/-5=x x=18/5 ok?

OpenStudy (anonymous):

yes